OpenStudy (anonymous):

Which point is NOT a solution of A. (0, 4) B. (-2, 3) C. (4, 4) D. (6, 0)

OpenStudy (anonymous):

we need the function/equation/formula

OpenStudy (anonymous):

x+2y≤8

OpenStudy (anonymous):

okay so try each answer. ill try A for you \[x + 2y \le 8\] \[0 + 2(4) \le 8\] \[8 \le 8\] which is true 8 is equal to 8 so were inserting the x and y values from the coordinate (x, y) and placing them into the equation where the letters x and y are. so you try b where (x, y) is (-2, 3) that is x = -2 and y = 3